Fermi National Laboratory
 

Accelerator Update for Jun 5, 2006 - June 6, 2006

 

Monday June 5

The day shift began with the Main Injector (MI), Antiproton Source (Pbar), Recycler, Switchyard (SY), NuMI, MiniBooNE, and the Tevatron (TeV) conducting studies, and with NuMI taking beam.

At 8:36 AM, Operations reported a LCW leak in a MI RF station (MIRF5). Operators turned off and bypassed the station. RF experts had it back on line an hour later.

Operations put the TeV into a dry squeeze at 8:45 AM for the first of five times over the next six and a half hours.

Operations prepared MiniBooNE for an access at 9:21 AM. The experimenters need to check their dewatering system.

Operations reported at 9:35 AM that the MI orbit shift was due to a bad power supply (I:QT101). Technicians replaced the supply and MI beam returned at 10:20 AM.

Operations reported 12:46 PM that the Pbar Accumulator had been collecting antiprotons (stacking) all morning.

At 12:46 PM, Operations reported that MiniBooNE was being held off due to a power supply problem (HV860). A power supply expert fixed the problem. Operations established beam to MiniBooNE at 12:59 PM.

Operations reported at 1:37 PM that MI experts began trying to send beam to Pbar and to NuMI at the same time. By 1:55 PM experts succeeded in establishing a mixed NuMI and stacking mode.

At 2:04 PM, Booster experts began their two-hour Linac-Booster energy match tune. This limited beam to all other downstream machines and cut short the mixed mode study.

Pbar experts began working on stacktail cooling at 2:23 PM.

Operators searcedh and secured the Meson enclosures M03, M04, and M05 at 3:22 PM. Interlock technicians began testing the safety system.

At 3:54 PM, the TeV power supply interlock loop tripped off, aborting the beam.

MI, Pbar, and NuMI experts resumed running mixed mode stacking studies at 4:05 PM.

At 4:08 PM, an expert investigated a problem with the TeV sector F2 dump.

NuMI reported at 5:22 PM that their BPM and auto-tune systems were working normally.

Operations reported at 5:25 PM that MI had finished their mixed mode studies. Experts had tried to run mixed mode without slip stacking, but it caused a trip. They will run with separate stacking and NuMI modules throughout the night.

Operations turned the TeV back on at 5:43 PM. Experts had found and replaced a blown fuse on the dump control chassis at F3.

Operations brought the TeV up to 980 GeV at 6:05 PM and then put the machine into a dry squeeze.

Tuesday June 6

The midnight shift began with TeV and Pbar experts conducting studies, and with MiniBooNE and NuMI taking beam.

At 12:09 AM, Operations reported that a water heater had tripped off on a ground fault indication for a TeV RF station (TRF1). Operator eventually bypassed this station.

At 3:39 AM, operators noticed an odd smell around a Booster RF station (BRF19). Later they noticed that BRF19 wasn’t operating normally. At 4:44 AM, they bypassed the station.

The Plans for Tuesday

The plan is to continue commissioning. The goals for today are to allow a brief access later this morning to various areas, and then to send antiprotons to the Recycler this afternoon. D0 will begin their sixteen-hour access on Wednesday at 7 AM.

Machine Reports

Linac

Okay

Booster

Experts only managed to get 0.02 percent change out of their energy match. They will continue on with commissioning from there.

Main Injector

Experts worked on slip stacking. They are working on a problem with BPMs.

Antiproton Source

Experts have the machine stacking at a rate of 2.7 milliamps per hour. They have phased in their stacktail system. Next they have to phase their Debuncher and their Accumulator core system. The Accumulator DCCT broke; an access is needed to repair this and, if there’s enough time, many other small jobs.

TeV

Experts continue to commission the machine. They have completed a 150 GeV feeddown scan, drift measurements, and aperture scans. Experts want to power the Tevatron Electron Lens (TEL2). They also want to access to work on TEL1 among other jobs.

Recycler

Experts are ready for antiprotons. They have been working with protons. An expert from the company that built the Pelletron resolved some of the problems. They ran the Pelletron all night and it didn’t trip. The cooling skid provided better cooling, but not what it was designed to provide. They have seen an electron beam make the round trip.

Switchyard 120

Experts tested power supplies. The A0 quadrupole magnet will be installed tomorrow during the D0 access.

MiniBooNE

Experts felt that they ran well.

NuMI

Experts reported that the mixed mode studies didn’t work as expected.

Collider Detector at Fermilab

CDF experts ran DAQ and trigger testing. They will access their hall today and on Wednesday.

D-Zero

Experts continue to investigate their noise problem. They will access their hall today and make a sixteen-hour access on Wednesday to make repairs.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 06/06/2006    email Fermilab